# Conformal online model aggregation

### v1

@@ 5 COMA under distribution shift
As explained in Section 1, conformal prediction offers a general approach to convert the output of any prediction algorithm into a prediction set. Although its validity hinges on the assumption of exchangeability, this assumption is frequently violated in numerous real-world applications. In the energy market, for instance, consumer behavior can undergo significant shifts in response to new legislation, major global events, and financial market fluctuations. In general, where the distribution of the data shift over time we cannot expect to obtain reliable results using standard conformal inference. In recent years, several works have been proposed to address this issue; some examples are Gibbs and Candès, (2021) and Angelopoulos et al., 2024a .
@@
$\alpha_{k}^{(1)}=\alpha,$ (9)
$\alpha_{k}^{(t)}=\alpha_{k}^{(t-1)}+\gamma(\alpha-\phi_{k}^{(t-1)}),\quad t\geq 2,$ (10)
where $\phi^{(t)}_{k}=\mathds{1}\{Y_{t}\notin\mathcal{C}_{k}^{(t)}(\alpha_{k}^{(t)})\}$ is the sequence of miscoverage events obtained by the $k$-th agent and setting the error to $\alpha_{k}^{(t)}$, while $\gamma>0$ represents a step size parameter. Possible improvements are proposed in Gibbs and Candès, (2023) and Zaffran et al., (2022), in particular, they tune the parameter $\gamma$ over time to make the procedure more flexible. A drawback arises from the fact that in certain situations, ACI can return either $\mathbb{R}$ or $\emptyset$ as a set, attributed to the circumstance that in some iterations, the parameter $\alpha^{(t)}$ can be smaller than 0 or greater than 1.
Another method is presented in Angelopoulos et al., 2024a , where instead of updating the error level $\alpha$, there is an update of the quantile of the distribution of scores. In this setting, we define $s^{(t)}:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$ as the conformal score function, which measures the accuracy of the prediction at time $t$. Actually, for different models one can use different scores (e.g. absolute or signed residuals), but for clarity of exposition, we will use the same $s(\cdot,\cdot)$ for all experts. The procedure updates the quantiles as follows:
$q_{k}^{(1)}=q,$
$q_{k}^{(t)}=q_{k}^{(t-1)}+\gamma(\phi_{k}^{(t-1)}-\alpha),\quad t\geq 2,$
where $\phi_{k}^{(t-1)}=\mathds{1}\{y^{(t)}\notin\mathcal{C}_{k}^{(t)}(q_{k}^{(t)})\}$ and the sets at each time are simply defined by
$\mathcal{C}_{k}^{(t)}=\left\{y\in\mathcal{Y}:\,s^{(t)}(x^{(t)},y)\leq q_{k}^{(t)}\right\}.$
Also in this case some improvements are proposed, such as an adaptive learning rate or the introduction of an error integrator. The proposal in Angelopoulos et al., 2024b is to set a decaying step size $\gamma^{(t)}$. In particular, they prove that if scores are bounded and $\gamma^{(t)}\propto t^{-1/2-\epsilon}$ for some $\epsilon\in(0,1/2)$, then the miscoverage coverage error is bounded as $\mathcal{O}(\frac{1}{(T^{1/2}-\epsilon)})$ in the adversarial setting, while in the i.i.d. the algorithm achieves a coverage guarantee.
@@ 5.1 Decentralized COMA under distribution shift
In the scenario described above, suppose that each agent operates independently by providing an interval at each iteration and, subsequently, the aggregator merges these intervals according to the vector $w^{(t)}$. This framework can be referred to as a decentralized dynamic merging. Under this setting, the advantage of the quantile tracking method is that it never returns sets of infinite dimensions. This implies that the loss function used can remain the size of the different sets, ensuring that the bound described in (8) is valid. In the case of using the described ACI approach, it becomes necessary to employ a function $g(\cdot)$ that is capable of returning a finite number when the miscoverage rate $\alpha_{k}^{(t)}$ is less than or equal to 0.
@@
We use the methods described in Section 3 on adaptive conformal inference in a real data set. We use $K=3$ different regression algorithms, and we employ quantile tracking with a decaying step size $\gamma^{(t)}\propto t^{-1/2-\epsilon}$. In particular, we use the ELEC2 (Harries,, 2003) data set that monitors electricity consumption and pricing in the states of New South Wales and Victoria in Australia, with data recorded every 30 minutes over a 2.5-year period from 1996 to 1999. For our experiment, we utilize four covariates: nswprice and vicprice, representing the electricity prices in each respective state, and nswdemand and vicdemand, denoting the usage demand in each state. The response variable is transfer, indicating the quantity of electricity transferred between the two states. We narrow our focus to a subset of the data, retaining only observations within the time range of 9:00 am to 12:00 pm to mitigate daily fluctuation effects, the same procedure is chosen in Barber et al., (2023). Additionally, we discard an initial time segment during which the transfer value remains constant. The models used correspond to three distinct linear models based on different covariates. The first model relies on the entire set of covariates at time $t$, the second model employs $y^{(t-1)}$ and $y^{(t-2)}$ as covariates (it corresponds to an AR(2) without constraints on the parameters), while the third model supplements the latter with the covariate nwsprice.
@@
Figure 2: The top row displays local coverage for adaptive conformal inference using quantile tracking for the three separate regression algorithms. The bottom row shows local coverage for our merged set using COMA with adaptive $\eta$ or $\eta=0.1$ (with and without randomization). In the bottom-left plot the lines of the randomized and non-randomized methods overlap as the weights concentrate after a few iterations onto a single model.
@@
$\textrm{localCov}^{(t)}:=\sum_{i=t-100+1}^{t+100}\mathds{1}\left\{y^{(i)}\in\mathcal{C}^{(i)}\left(q_{k}^{(i)}\right)\right\},$ (11)
where $\mathcal{C}_{k}^{(t)}(q_{k}^{(t)})$ is the set obtained using the $k$-th regression algorithms with the quantile fixed at $q_{k}^{(t)}$. The local coverage is also used to measure the performance of our merging procedure. In particular, we compare the use of an adaptive learning parameter and a small fixed $\eta=0.1$. As can be seen in Figure 2, the COMA procedure with adaptive $\eta^{(t)}$ obtains a coverage that is similar to the prespecified level $1-\alpha$. This is because the weights concentrate rapidly on one model over time; in particular, this model is not always the same, as is possible from Figure 3. However, in both cases, the coverage is similar to the pre-specified level $1-\alpha$.
Figure 3: Weights (first row) and losses, $\ell^{(t)}$, (second row) obtained by the various model for the decentralized COMA algorithm applied in ELEC2 dataset. The series of the losses have been smoothed using a moving average $(100,\frac{1}{100})$. The loss of the weighted majority vote stabilizes around the two best models after few iterations. The strategy with an adaptive $\eta^{(t)}$ is more aggressive than the one with $\eta^{(t)}=\eta$.
@@ 5.2 Adaptive conformal inference directly applied on dynamic merging
$\mathcal{C}^{(t)}:=\left\{y\in\mathbb{R}:\sum_{k=1}^{K}w_{k}^{(t)}\mathds{1}\left\{y\in\mathcal{C}_{k}^{(t)}\left(\alpha^{(t)}\right)\right\}>1/2\right\},$
@@
Figure 4: Local level coverage, weights assumed during the iterations and losses for direct ACI for COMA with adaptive $\eta$ (top) and $\eta=0.1$ (bottom) applied in ELEC2 dataset. The series of the losses have been smoothed using a moving average $(100,\frac{1}{100})$. In both cases, the local level coverage oscillates around the level $1-\alpha$.
@@ 6 Summary
Our work presents a method for aggregating conformal prediction intervals derived from different algorithms. Specifically, a method based on majority voting is proposed, where the weight of each individual agent is updated based on its past performance. In some cases, when there is an algorithm that outperforms the others, the weight is assigned to the better-performing model, resulting in model selection. Particularly, in the case of i.i.d. data, it is very common for there to exist a better model that outperforms the others in terms of size of the outputted sets.
We also tested our method in cases where the data underwent a distribution shift, along with quantile tracking (Angelopoulos et al., 2024a, ; Angelopoulos et al., 2024b, ) and adaptive conformal inference (Gibbs and Candès,, 2021) methods. Specifically, two contexts are explored: in the first case, each model operates independently, while in the second case, the aggregator acts above the models, and adaptive conformal inference operates directly on it.
@@ Appendix A Additional results from Section 4
Figure 7: Series of values taken by the learning parameter $\eta^{(t)}$ over time. The plot on the right corresponds to the plot on the left, zoomed in on the interval 0 to 10.
Appendix B Additional simulations on Adaptive Conformal Inference

### v2

@@ 5 COMA under distribution shift
As described in Section 1, conformal prediction offers a general approach to convert the output of any prediction algorithm into a prediction set. Although its validity hinges on the assumption of exchangeability, this assumption is frequently violated in numerous real-world applications. In the energy market, for instance, consumer behavior can undergo significant shifts in response to new legislation, major global events, and financial market fluctuations. In general, where the distribution of the data shift over time we cannot expect to obtain reliable results using standard conformal inference. In recent years, several works have been proposed to address this issue; some examples are Gibbs and Candès, (2021) and Angelopoulos et al., 2024a . For the introduction of the methods, in order to avoid confusion, the discussion refers to situations involving the use of just one algorithm. The case involving multiple experts is addressed in Section 5.1 and Section 5.2.
@@
$\alpha^{(1)}=\alpha,$ (10)
$\alpha^{(t)}=\alpha^{(t-1)}+\gamma(\alpha-\phi^{(t-1)}),\quad t\geq 2,$ (11)
where $\phi^{(t)}=\mathds{1}\{Y_{t}\notin\mathcal{C}^{(t)}(\alpha^{(t)})\}$ is the sequence of miscoverage events setting the error to $\alpha^{(t)}$, while $\gamma>0$ represents a step size parameter. Possible improvements are proposed in Gibbs and Candès, (2023) and Zaffran et al., (2022), in particular, their proposals tune the parameter $\gamma$ over time using an online learning procedure. A drawback arises from the fact that in certain situations, ACI can return either $\mathbb{R}$ or $\emptyset$ as a set, attributed to the circumstance that in some iterations, the parameter $\alpha^{(t)}$ can be smaller than 0 or greater than 1.
Another method is presented in Angelopoulos et al., 2024a , where instead of updating the error level $\alpha$, there is an update of the quantile of the distribution of the scores. In this setting, we define $s^{(t)}:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$ as the conformal score function, which measures the accuracy of the prediction at time $t$. Actually, if different models are involved one can use different scores (e.g. absolute or signed residuals), but in the following we will assume the same $s(\cdot,\cdot)$ for all algorithms. The procedure updates the quantiles as follows:
$q^{(1)}=q,$
$q^{(t)}=q^{(t-1)}+\gamma(\phi^{(t-1)}-\alpha),\quad t\geq 2,$
where $\phi^{(t)}=\mathds{1}\{y^{(t)}\notin\mathcal{C}^{(t)}(q^{(t)})\}$ and the sets at each time are simply defined by
$\mathcal{C}^{(t)}(q^{(t)})=\left\{y\in\mathcal{Y}:\,s^{(t)}(x^{(t)},y)\leq q^{(t)}\right\}.$
As noted in Angelopoulos et al., 2024a , both ACI and quantile tracking, can be recasted to an online descent algorithm with respect to the pinball loss (Koenker,, 2005). In addition, both methods achieve a long-run coverage in time, defined as
$\frac{1}{T}\sum_{t=1}^{T}\phi^{(t)}=\alpha+o(1),$ (12)
under few or no assumptions on $\{z^{(t)}\}_{t=1}^{T}$ and $\{s^{(t)}\}_{t=1}^{T}$, where $o(1)$ is quantity that tends to zero as $T\to\infty$. In particular, (15) holds deterministically and is different from the probabilistic assumption in (1) valid in an i.i.d. setting.
Also in the case of the quantile tracking some improvements are proposed, such as an adaptive learning rate or the introduction of an error integrator. The proposal in Angelopoulos et al., 2024b is to set a decaying step size $\gamma^{(t)}$. In particular, they prove that if scores are bounded and $\gamma^{(t)}\propto t^{-1/2-\epsilon}$ for some $\epsilon\in(0,1/2)$, then the miscoverage coverage error is bounded as $O(\frac{1}{T^{1/2-\epsilon}})$ in the adversarial setting, while in the i.i.d. the algorithm achieves a coverage guarantee.
@@ 5.1 Decentralized COMA under distribution shift
In the scenario described above, suppose that each agent operates independently by providing an interval at each iteration and, subsequently, the aggregator merges these intervals according to the vector $w^{(t)}$. This framework can be referred to as a decentralized dynamic merging. Under this setting, the advantage of the quantile tracking method is that it never returns sets of infinite dimensions. This implies that the loss function used can remain the size of the different sets, ensuring that the bound described in (9) is valid. In the case of using the described ACI approach, it becomes necessary to employ a function $g(\cdot)$ that is capable of returning a finite number when the miscoverage rate $\alpha_{k}^{(t)}$ is smaller than 0.
Before proceeding with the rest of the section, we introduce the following proposition, which holds under the following assumptions:
  1. 1.
    Negative total elementwise empirical correlation between $\{\phi^{(t)}\}_{t=1}^{T}$ and $\{w^{(t)}\}_{t=1}^{T}$:
    $(T^{-1})\sum_{t=1}^{T}\sum_{k=1}^{K}(\phi_{k}^{(t)}-\bar{\phi}_{k})(w_{k}^{(t)}-\bar{w}_{k})\leq 0,$ (13)
    where $\bar{\phi}_{k}:=(T^{-1})\sum_{t=1}^{T}\phi_{k}^{(t)}$ and $\bar{w}_{k}:=(T^{-1})\sum_{t=1}^{T}w_{k}^{(t)}$.
  2. 2.
    No randomization: the value $u^{(t)}$ in (3) equals 0, in symbols,
    $u^{(t)}=0,\quad t=1,\dots,T.$ (14)
Proposition 5.1.
Let $\{z^{(t)}=(x^{(t)},y^{(t)})\}_{t=1}^{T}$ be a sequence of data points, $\{w^{(t)}=(w_{1}^{(t)},\dots,w_{k}^{(t)})\}_{t=1}^{T}$ be a sequence of weight vectors and $\{\phi_{k}^{(t)}\}_{t=1}^{T}$ be the sequence of miscoverage events for the $k$-th algorithm, $k=1,\dots,K$. Under the assumptions in (13) and (14), then the decentralized dynamic merging in conjunction with adaptive conformal inference defined in (10) and (11) satisfies
$\frac{1}{T}\sum_{t=1}^{T}\mathds{1}\{y^{(t)}\notin\mathcal{C}^{(t)}\}\leq 2\alpha+2\frac{1-\alpha+\gamma}{T\gamma},$ (15)
for all $\alpha\in[0,1/2]$.
The proof is given in Appendix C, and similar results can be obtained using the quantile tracking method under the assumption of bounded scores (see Proposition 1 in Angelopoulos et al., 2024a and Theorem 1 in Angelopoulos et al., 2024b ). Randomization is not included because the UMI inequality is inapplicable in deterministic contexts however, the results using randomization appear to be reasonably decent also in this setting. The result (15) expresses a long-run coverage guarantee of level $1-2\alpha$ under arbitrary distribution if certain conditions regarding the relationship between the weights and the miscoverage events are fulfilled. Such conditions are challenging to verify in reality; however, empirically, the method appears to yield satisfactory results in real-world applications (see Appendix B for other experiments).
@@
We now use the methods described in Section 3 in a real data set. We define $K=3$ different regression algorithms, and we employ quantile tracking with a decaying step size $\gamma^{(t)}\propto t^{-1/2-\epsilon}$. In particular, we use the ELEC2 (Harries,, 2003) data set that monitors electricity consumption and pricing in the states of New South Wales and Victoria in Australia, with data recorded every 30 minutes over a 2.5-year period from 1996 to 1999. For our experiment, we utilize four covariates: nswprice and vicprice, representing the electricity prices in each respective state, and nswdemand and vicdemand, denoting the usage demand in each state. The response variable is transfer, indicating the quantity of electricity transferred between the two states. We narrow our focus to a subset of the data, retaining only observations within the time range of 9:00 am to 12:00 pm to mitigate daily fluctuation effects, the same procedure is chosen in Barber et al., (2023). Additionally, we discard an initial time segment during which the transfer value remains constant. The models used correspond to three distinct linear models based on different covariates. The first model relies on the entire set of covariates at time $t$, the second model employs $y^{(t-1)}$ and $y^{(t-2)}$ as covariates (it corresponds to an AR(2) without constraints on the parameters), while the third model supplements the latter with the covariate nwsprice.
@@
Figure 2: The top row displays local coverage for adaptive conformal inference using quantile tracking for the three regression algorithms described in Section 5.1. The bottom row shows local coverage for our merged set using COMA with adaptive $\eta$ or $\eta=0.1$ (with and without randomization). In the first row the dashed lines represent the level $1-\alpha$, while in the second column dashed lines represent the level $1-2\alpha$. In the bottom-left plot the lines of the randomized and non-randomized methods overlap as the weights concentrate after a few iterations onto a single model.
@@
$\textrm{localCov}^{(t)}:=\sum_{i=t-100+1}^{t+100}\mathds{1}\left\{y^{(i)}\in\mathcal{C}_{k}^{(i)}\left(q_{k}^{(i)}\right)\right\},$ (16)
where $\mathcal{C}_{k}^{(t)}(q_{k}^{(t)})$ is the set obtained by the $k$-th regression algorithms with the quantile fixed at $q_{k}^{(t)}$. The local coverage is also used to measure the performance of our merging procedure. In particular, we compare the use of an adaptive learning parameter and a small fixed $\eta=0.1$. As can be seen in Figure 2, the COMA procedure with adaptive $\eta^{(t)}$ obtains a coverage that is similar to the prespecified level $1-\alpha$. This is because the weights concentrate rapidly on one model over time; in particular, this model is not always the same, as is possible from Figure 3. However, in both cases, the coverage is similar to the pre-specified level $1-\alpha$.
Figure 3: Weights (first row) and losses, $\ell^{(t)}$, (second row) obtained by the various model for the decentralized COMA algorithm applied in ELEC2 dataset (Section 5.1). The series of the losses have been smoothed using a moving average $(100,\frac{1}{100})$. The loss of the weighted majority vote stabilizes around the two best models after few iterations. The strategy with an adaptive $\eta^{(t)}$ is more aggressive than the one with $\eta^{(t)}=\eta$.
@@ 5.2 Adaptive conformal inference directly applied on dynamic merging
$\mathcal{C}^{(t)}:=\left\{y\in\mathcal{Y}:\sum_{k=1}^{K}w_{k}^{(t)}\mathds{1}\left\{y\in\mathcal{C}_{k}^{(t)}\left(\alpha^{(t)}\right)\right\}>1/2\right\},$
@@
Figure 4: Local level coverage (first column), weights assumed during the iterations (second column) and losses (third column) for direct ACI for COMA with adaptive $\eta$ (top) and $\eta=0.1$ (bottom) applied in ELEC2 dataset (Section 5.2). The dashed line in the first column represents the level $1-\alpha$. The series of the losses have been smoothed using a moving average $(100,\frac{1}{100})$. In both cases, the local level coverage oscillates around the level $1-\alpha$.
@@ 6 Summary
Our work presents a method for aggregating conformal prediction intervals derived from different algorithms. Specifically, a method based on majority voting is proposed, where the weight of each individual agent is updated based on its past performance. In some cases, when there is an algorithm that outperforms the others, the weight is assigned to the better-performing model, resulting in model selection. Particularly, in the case of i.i.d. data, it is very common for there to exist a best model that outperforms the others in terms of size of the outputted sets. It is feasible to establish an upper bound for the loss incurred by our method relative to the loss of the best-performing algorithm, translating into a form of regret for our method.
We also tested our method in cases where the data underwent a distribution shift, along with quantile tracking (Angelopoulos et al., 2024a, ; Angelopoulos et al., 2024b, ) and adaptive conformal inference (Gibbs and Candès,, 2021) methods. Specifically, two contexts are explored: in the first case, each model operates independently, while in the second case, the aggregator acts above the models, and adaptive conformal inference operates directly on it. Furthermore, even when the assumptions are not fully satisfied, the empirical results appear to meet the coverage expectations. Further research is required to assess whether our approaches can protect against a broader range of real-world situations.
@@ Appendix A Additional results from Section 4
Figure 7: Series of values taken by the learning parameter $\eta^{(t)}$ over time. The plot on the right corresponds to the plot on the left, zoomed in on the interval 0 to 10. The red dashed lines and the green dashed line represent the level 1 and 0.01. Solid lines denote the trajectory of the parameter $\eta^{(t)}$.
Appendix B Further experiments
B.1 Additional simulations on Adaptive Conformal Inference
@@
B.2 Amazon data
Inspired by Angelopoulos et al., 2024a , we analyze a historical series comprising the single stock open prices of Amazon from January 1, 2006, to December 31, 2014. The compared models consist of six different autoregressive models incorporating temporal lags from 1 to 6, trained on the logarithmic scale of the opening price, while the forecasts are obtained on the original scale. The method used to obtain the sets is the quantile tracking method described in Section 5 with $\alpha=0.1$, a similar analysis is performed in Angelopoulos et al., 2024a with an autoregressive model with 3 lags. The parameter $\gamma$ is updated as $\gamma^{(t)}=0.1B^{(t)}$, where $B^{(t)}$ is the maximum observed score in a trailing window of 100 iterations and the models are re-trained at every time step. The weights are updated using an adaptive $\eta$ or fixing $\eta$ to the level 0.01, this small value of $\eta$ allows the mixing of the different intervals during iterations.
Figure 9: Local level coverage obtained in a window of 100 data points (first column), weights assumed during the different iterations (second column), empirical sum over time of the quantity $\sum_{k=1}^{K}\phi_{k}^{(t)}w_{k}^{(t)}-\sum_{k=1}^{K}\phi_{k}^{(t)}\sum_{k=1}^{K}w_{k}^{(t)}$ obtained using an increasing number of observations (third column) and series of the original stock with corresponding prediction intervals (fourth column). The black dashed lines in the first column represent the level $1-2\alpha$. The result corresponding to the strategy with an adaptive $\eta$ are presented in the first row while the result corresponding to the strategy with $\eta=0.01$ are presented in the second row.
As shown in Figure 9, the local leverage coverage exceeds the level $1-2\alpha$ and approximates $1-\alpha$ in both instances. The approach using an adaptive $\eta$ aggressively allocates weights and it frequently performs model selection. Nonetheless, in each scenario, the AR(1) model outperforms others in the initial phase, whereas the AR(4) model appears more effective in the latter phase. The empirical sum of the quantity $\sum_{k=1}^{K}\phi_{k}^{(t)}w_{k}^{(t)}-\sum_{k=1}^{K}\phi_{k}^{(t)}\sum_{k=1}^{K}w_{k}^{(t)}$ it is around zero during all iterations in both cases.
B.3 Lasso regression with different tuning parameters
We create a synthetic dataset regarding a high-dimensional linear model in this way. We first simulate a matrix $X$ of dimension $3100\times 100$ from a multivariate normal with variance $I_{100}$, where $I_{a}$ denote an identity matrix of dimension $a\times a$. The outcome vector is defined as follows:
$y^{(t)}=\beta^{\mathrm{T}}x^{(t)}+0.3y^{(t-1)}+\varepsilon^{(t)}-0.3\varepsilon^{(t-1)},$
where $\beta$ is a coefficient vector with non-zero entries only in the first ten positions, $x^{(t)}$ represents the $t$-th row vector of $X$, and $\varepsilon^{(t)}$ is sampled from a standard Gaussian distribution. The data generation process follows an ARMA(1,1) model (see Chapter 2.6 in Tsay, (2005) for an introduction), augmented by a linear component that depends on the covariates in $X$. Here, the employed models are lasso regressions, each with varying penalty parameters $\lambda$. In particular, the $\lambda$ values are logarithmically equispaced in the interval $[-4,3]$. As the data are not exchangeable, the quantile tracking approach is utilized along with COMA to obtain prediction intervals. Throughout all iterations, the parameter $\gamma$ is set to 2.
In Figure 10, the best $\lambda$ oscillates between $\lambda_{4}$ and $\lambda_{5}$ for all iterations. The local level coverage is between the level $1-2\alpha$ and $1-\alpha$ in both cases. In the last column, it is possible to see that the difference between $\sum_{k=1}^{K}\phi_{k}^{(t)}w_{k}^{(t)}$ and $\sum_{k=1}^{K}\phi_{k}^{(t)}\sum_{k=1}^{K}w_{k}^{(t)}$ is around the zero level and decreases over time.
Figure 10: Local level coverage obtained in a window of 100 data points (first column), weights assumed during the different iterations (second column) and empirical sum over time of the quantity $\sum_{k=1}^{K}\phi_{k}^{(t)}w_{k}^{(t)}-\sum_{k=1}^{K}\phi_{k}^{(t)}\sum_{k=1}^{K}w_{k}^{(t)}$ obtained using an increasing number of observations (third column). The black dashed lines in the first column represent the level $1-2\alpha$. The result corresponding to the strategy with an adaptive $\eta$ are presented in the first row while the result corresponding to the strategy with $\eta=0.01$ are presented in the second row.
B.4 Additional simulations on the negative correlation assumption
In this section, we investigate, the assumption of negative correlation via simulation. Data are generated as follows:
$y^{(t)}=10\sin(\pi x_{1}^{(t)})+10\exp\{x_{2}^{(t)}\}-4x_{3}^{(t)}+0x_{4}^{(t)}+\varepsilon^{(t)},\quad t=1,\dots,300,$ (17)
where $x_{j}^{(t)}$, $j=1,\dots,4$, are generated from independent uniform random variable in (0,1) while $\varepsilon^{(t)}\sim\mathcal{N}(0,1)$. The burn-in period for training is equal to 100 observations and four different models are used: random forest, linear model, lasso, and neural net. The number of replications is equal to 2000 and all the results in Figure 11 are averaged over the number of replications.
Figure 11: Empirical correlation between $\phi_{k}^{(t)}$ and $w_{k}^{(t)}$ at every time step (first columns), empirical average of the weights (second column) and empirical calculation of the value $\sum_{k=1}^{K}(\mathbb{E}[W_{k}^{(t)}\phi_{k}^{(t)}]-\mathbb{E}[W_{k}^{(t)}]\mathbb{E}[\phi_{k}^{(t)}])$ (third column).
From Figure (11), it seems that the assumption of negative correlation is not always met for all iterations for each model, however, the empirical difference between $\sum_{k=1}^{K}\mathbb{E}[W_{k}^{(t)}\phi_{k}^{(t)}]$ and $\sum_{k=1}^{K}\mathbb{E}[W_{k}^{(t)}]\mathbb{E}[\phi_{k}^{(t)}]$ does not appear to be high. In this case, the neural net appears to be the best model, followed by the random forest. Indeed, this can be inferred from the non-linear relationship between the covariates and the response in (17).
B.5 Classification task in an online setting
In this section, we tackle the problem of classification within an online framework. A simulated dataset with eight distinct classes were generated, each class is originated from a bivariate normal distribution with mean vectors of $(i,i)$ and randomly generated covariance matrices for $i=1,\dots,8$. Additionally, two covariates were introduced to induce noise. We generated 250 data points per class, and treated the classification process as if the data points were arriving in a sequential manner. Specifically, following a burn-in phase of 1000 iterations, we retrained four different classifiers (LDA, QDA, Neural Net, Random Forest) at each step, and the weights for these algorithms were adjusted using the COMA method.
Figure 12: Local level coverage obtained in a window of 100 data points (first column), weights assumed during the different iterations (second column), empirical sum over time of the quantity $\sum_{k=1}^{K}\phi_{k}^{(t)}w_{k}^{(t)}-\sum_{k=1}^{K}\phi_{k}^{(t)}\sum_{k=1}^{K}w_{k}^{(t)}$ obtained using an increasing number of observations (third column) and series of the original stock with corresponding prediction intervals (fourth column). The black dashed lines in the first column represent the level $1-2\alpha$. The result corresponding to the strategy with an adaptive $\eta$ are presented in the first row while the result corresponding to the strategy with $\eta=0.01$ are presented in the second row.
The results are presented in Figure 12, where it is possible to note that the miscoverage lies between the level $1-\alpha$ and $1-2\alpha$. In this situation, the strategy with an adaptive $\eta$ gives very different weights if compared to the strategy with a $\eta=0.01$. In the last column, the sum of covariances is around zero for all iterations.
Appendix C Omitted proofs
Proof of Theorem 2.1.
By direct calculation, we can write
$\begin{split}\mathbb{P}\left(Y^{(t)}\notin\mathcal{C}^{(t)}\right)&=\mathbb{P}\left(\sum_{k=1}^{K}W_{k}^{(t)}\phi_{k}^{(t)}\geq\frac{1}{2}-U^{(t)}/2\right)\stackrel{{\scriptstyle(i)}}{{\leq}}2\mathbb{E}\left[\sum_{k=1}^{K}W_{k}^{(t)}\phi_{k}^{(t)}\right]=2\sum_{k=1}^{K}\mathbb{E}\left[W_{k}^{(t)}\phi_{k}^{(t)}\right]\\ \end{split}$
where $(i)$ is valid due to the uniformly-randomized Markov’s inequality described in Ramdas and Manole, (2023).
The first part of the theorem is a direct consequence of Hölder inequality. Indeed, due to the fact that $||W||_{1}=1$, we notice that
$\mathbb{E}\left[\sum_{k=1}^{K}\phi_{k}^{(t)}W_{k}^{(t)}\right]\leq\mathbb{E}\left[||\phi^{(t)}||_{\infty}||W^{(t)}||_{1}\right]=\mathbb{E}\left[||\phi^{(t)}||_{\infty}\right],$
where $\phi^{(t)}=(\phi_{1}^{(t)},\dots,\phi_{K}^{(t)})$.
The second part can be proved as follows: in the case of conditional coverage (case 1.), then it holds that
$\begin{split}\mathbb{P}\left(Y^{(t)}\notin\mathcal{C}^{(t)}\right)&\leq 2\sum_{k=1}^{K}\mathbb{E}\left[W_{k}^{(t)}\phi_{k}^{(t)}\right]=2\sum_{k=1}^{K}\mathbb{E}\left[\mathbb{E}\left[W_{k}^{(t)}\phi_{k}^{(t)}\mid\{Z^{(i)}\}_{i=1}^{T}\right]\right]\\ &=2\sum_{k=1}^{K}\mathbb{E}\left[W_{k}^{(t)}\mathbb{E}\left[\phi_{k}^{(t)}\mid\{Z^{(i)}\}_{i=1}^{T}\right]\right]\leq 2\alpha.\end{split}$
In the case of negative total elementwise correlation (case 2.) it holds that $\sum_{k=1}^{K}\mathbb{E}[W_{k}^{(t)}\phi_{k}^{(t)}]\leq\sum_{k=1}^{K}\mathbb{E}[W_{k}^{(t)}]\mathbb{E}[\phi_{k}^{(t)}]$ and under marginal coverage, we have that
$\mathbb{P}\left(Y^{(t)}\notin\mathcal{C}^{(t)}\right)\leq 2\sum_{k=1}^{K}\mathbb{E}\left[W_{k}^{(t)}\phi_{k}^{(t)}\right]\leq 2\sum_{k=1}^{K}\mathbb{E}[W_{k}^{(t)}]\,\mathbb{E}[\phi_{k}^{(t)}]\leq 2\alpha\sum_{k=1}^{K}\mathbb{E}[W_{k}^{(t)}]=2\alpha.$
This concludes the theorem. ∎
Proof of Proposition 5.1.
We remark that, if $u^{(t)}=0$, then $y^{(t)}$ is not contained in $\mathcal{C}^{(t)}$ if and only if $\sum_{k=1}^{K}w_{k}^{(t)}\mathds{1}\{y^{(t)}\notin\mathcal{C}_{k}^{(t)}\}\geq 1/2$. Then, since $\phi_{k}^{(t)}:=\mathds{1}\{y^{(t)}\notin\mathcal{C}_{k}^{(t)}\}$, we can write
$\begin{split}\sum_{t=1}^{T}\mathds{1}\{y^{(t)}\notin\mathcal{C}^{(t)}\}&=\sum_{t=1}^{T}\mathds{1}\left\{\sum_{k=1}^{K}w_{k}^{(t)}\mathds{1}\{y^{(t)}\notin\mathcal{C}_{k}^{(t)}\}\geq\frac{1}{2}\right\}\\ &\stackrel{{\scriptstyle(i)}}{{\leq}}\sum_{t=1}^{T}2\sum_{k=1}^{K}w_{k}^{(t)}\mathds{1}\{y^{(t)}\notin\mathcal{C}_{k}^{(t)}\},\\ &=2\sum_{k=1}^{K}\sum_{t=1}^{T}w_{k}^{(t)}\phi_{k}^{(t)}\\ &\stackrel{{\scriptstyle(ii)}}{{\leq}}2\sum_{k=1}^{K}\frac{1}{T}\left(\sum_{t=1}^{T}w_{k}^{(t)}\right)\left(\sum_{t=1}^{T}\phi_{k}^{(t)}\right)\\ &\stackrel{{\scriptstyle(iii)}}{{\leq}}2\sum_{k=1}^{K}\frac{1}{T}\left(\sum_{t=1}^{T}w_{k}^{(t)}\right)\left(T\alpha+\frac{1-\alpha+\gamma}{\gamma}\right)\\ &=2\left(\alpha+\frac{1-\alpha+\gamma}{T\gamma}\right)\sum_{k=1}^{K}\sum_{t=1}^{T}w_{k}^{(t)}\\ &=2T\alpha+2\frac{1-\alpha+\gamma}{\gamma},\\ \end{split}$
where $(i)$ holds due to the fact that $\mathds{1}\{x\geq 1\}\leq x,$ if $x\geq 0$ and $(ii)$ is due to assumption (13), in fact if the assumption holds, then $T\sum_{t=1}^{T}w_{k}^{(t)}\phi_{k}^{(t)}-\sum_{t=1}^{T}\phi_{k}^{(t)}\sum_{t=1}^{T}w_{k}^{(t)}\leq 0$. Proposition 4.1 in Gibbs and Candès, (2021) along with $\alpha\in[0,1/2]$ implies $(iii)$. Dividing both sides by $T$ gives the result. ∎